Parallel beam of light parallel to the principal axis of a lens falls on a lens of focal length 50 cm. After passing through the first lens they pass through another coaxially mounted convex lens of focal length 50 cm placed 150 cm from the first lens. After passing through the second lens, the rays converge at a distance of ________________ from the first lens.

Grade:11

2 Answers

pankaj kumar raman
12 years ago

after passing from 1st lens

it will form image at 50 cm from 1st lens

i.e. 100 cm from 2nd lens

hence for 2nd lens the obj. dist.(ray here) will be 100

applying lens formula

the the image dist. will be 100 cm

 

hope yhat my answer would be correct

tanmay seth
12 years ago

simple,

since light beam is parallel it will converge at focus i.e at 50cm from 1st lense. secondly, the other lense is 150 cm away from 1st lense hence the image formed from the first lense will serve as object for second lense,so it is 150 - 50 = 100 cm away from second lense.Now we can use lense formula 1/v -1/u =1/f , u= -50cm ,f= 50cm v= ? ,hence

1/v - (1/-100) = 1/50 ----> 1/v= 1/50-1/100= 1/100 i.e v= 100 cm